OpenStudy (anonymous):

Solve for x -(1/7) = x + (1 2/3) I know you need to "balance" it, so you would subtract 1 and 2/3 from each side. But i keep gettinng confused during that so if someone could please walk me through it I would love you forever.

OpenStudy (anonymous):

x = (-29/7)

OpenStudy (alekos):

if we subtract 1 2/3 form each side we get -1/7 - 1 2/3 = x or x = -1/7 - 1 2/3 = -1/7 - 5/3 so now we place the rhs on a common denominator of 21 x = (-3 -35)/21 x = -38/21

OpenStudy (anonymous):

thank you so mtcg